OpenStudy (anonymous):

(-2,-7) & (-6,-4) slope intercept form please.

OpenStudy (anonymous):

they explained before how to do it.. try it.. step by step

OpenStudy (anonymous):

ok folks. Here is what I came up with, please verify. y=(3/4)x - 13/4

OpenStudy (anonymous):

work out with se signs

OpenStudy (anonymous):

with the signs

OpenStudy (anonymous):

what do you mean se signs

OpenStudy (anonymous):

(y-(-7))=3/4(x-(-2))

OpenStudy (anonymous):

y+7=3/4(x+2)

OpenStudy (anonymous):

y+7=3/4x+2 y=3/4x+11/2

OpenStudy (anonymous):

? slope should be -3/4

OpenStudy (anonymous):

slope is +3/4

OpenStudy (anonymous):

{-4-(-7)} / {-6-(-2)} = 3/(-4) = -3/4

OpenStudy (anonymous):

so the slopw intercept form will be y=-3/4x-11/2

OpenStudy (anonymous):

nop... is what i said

OpenStudy (anonymous):

I tried to post this but had problems. The answer was -3/4x - 17/2 so what did we do wrong?

OpenStudy (anonymous):

is good

OpenStudy (anonymous):

it is like that

OpenStudy (anonymous):

i posted on your new one ...